# S2AFormer: Strip Self-Attention for Efficient Vision Transformer

### v1

@@ III-C Complexity Analysis
ResNet-18 [63] - - 32.9
EMO-1M [13] 5 23 34.2
PVT-T [25] 17 33 35.7
ResNet-50 [63] 29 46 36.7
S2AFormer-mini 6 23 36.7
PoolFormer-S12 [30] 16 31 37.2
PVTv2-B0 [16] 8 25 37.2
EMO-2M [13] 6 24 37.3
CASViT-XS [61] 7 24 37.1
S2AFormer-T 7 25 38.0
FastViT-SA12 [14] 14 29 38.0
ResNet-101 [63] 48 65 38.8
EfficientFormer-L1 [43] 16 28 38.9
ResNeXt-101-32x4d [80] 47 65 39.7
EdgeViT-XXS [12] 8 24 39.7
S2AFormer-XS 8 25 39.2
ResNeXt-101-64x4d [80] 86 104 40.2
EMO-5M [13] 9 26 40.4
PVT-S [25] 28 45 39.8
RepViT-M1.1 [31] - - 40.6
PoolFormer-S24 [30] 25 39 40.3
S2AFormer-S 12 28 40.8
PoolFormer-S36 [30] 35 48 42.0
EfficientFormerV2-S2 [64] 16 28 42.4
PVTv2-B1 [16] 18 34 42.5
PoolFormer-M48 [30] 77 82 42.7
InceptionNeXt-T [17] 28 44 43.1
RepViT-M1.5 [31] - - 43.6
S2AFormer-M 26 43 43.7
TABLE V: Throughput comparison across GPU, CPU, and ONNX platforms. GPU: NVIDIA H100 NVL, CPU: Intel(R) Xeon(R) Gold 6426Y.
Model #P.$\downarrow$ Top-1$\uparrow$ Throughput (images/s)
[…]
EMO-1M [13] 1.3 71.5 4674 45.57 118.72
EfficientViT-M3 [11] 6.9 73.4 4355 35.34 169.35
EfficientViT-M4 [11] 8.8 74.3 4289 34.34 159.51
EMO-2M [13] 2.3 75.1 3359 41.20 83.79
PVT-T [25] 13.2 75.1 5033 61.01 117.63
S2AFormer-mini 5.0 75.1 6330 58.21 118.26
FastViT-T8 [14] 3.6 75.6 7615 50.76 129.27
EfficientFormerV2-S0 [64] 3.5 75.7 520 44.57 136.32
EdgeViT-XXS [12] 4.1 74.4 3614 50.82 134.79
EfficientViT-M5 [11] 12.4 77.1 3843 28.48 126.28
EdgeViT-XS [12] 6.7 77.5 2872 35.14 106.10
S2AFormer-T 5.8 77.7 4768 41.75 112.89
EMO-5M [13] 5.1 78.4 2370 38.18 62.52
PVTv2-B1 [16] 14.0 78.7 3930 79.51 37.96
EMO-6M [13] 6.1 79.0 2102 35.90 57.00
EfficientFormerV2-S1 [64] 6.1 79.0 472 37.89 104.35
EfficientFormer-L1 [43] 12.3 79.2 3572 66.93 88.72
S2AFormer-XS 6.5 78.9 4185 34.09 71.63
FastViT-T12 [14] 6.8 79.1 5277 42.06 88.31
FastViT-S12 [14] 8.8 79.8 4952 39.41 68.09
FastViT-SA12 [14] 10.9 80.6 4812 38.48 50.12
PVT-S [25] 24.5 79.8 2736 43.50 53.01
EdgeViT-S [12] 11.1 81.0 1951 23.15 58.32
S2AFormer-S 10.7 80.8 2644 25.78 55.84
PVT-M [25] 44.2 81.2 1717 25.96 36.89
EfficientFormerV2-S2 [64] 12.6 81.6 256 25.08 46.94
PVT-L [25] 61.4 81.7 587 18.84 27.77
PVTv2-B2 [16] 25.4 82.0 2218 40.20 40.83
EfficientFormer-L3 [43] 31.3 82.4 1550 35.96 25.31
S2AFormer-M 24.9 82.3 1652 15.50 27.15
TABLE VI: Comparison of object detection and instance segmentation performance (%) on the COCO val2017 dataset [81]. FLOPs are tested on images of size $800\times 1280$. Our results, highlighted in bold, demonstrate superior performance with comparative computational overhead.
[…]
S2AFormer-mini 12/22 159/177 33.4 53.2 34.9 19.9 36.3 44.5 33.4 55.4 35.2 31.7 52.5 33.3
ResNet-18 [63] 21/31 -/- 31.8 49.6 33.6 16.3 34.3 43.2 34.0 54.0 36.7 31.2 51.0 32.7
EfficientViT-M4 [11] 9/- -/- 32.7 52.2 34.1 17.6 35.3 46.0 - - - - - -
PoolFormer-S12 [30] 22/32 -/- 36.2 56.2 38.2 20.8 39.1 48.0 37.3 59.0 40.1 34.6 55.8 36.9
CAS-ViT-XS [61] 12/23 162/181 36.5 56.3 38.9 21.8 39.9 48.4 37.4 59.1 40.4 34.9 56.2 37.0
S2AFormer-T 12/23 164/182 36.7 57.0 39.1 21.1 39.7 48.6 37.6 59.8 40.6 35.4 57.2 37.6
MF-508M [56] 8/- 168/- 38.0 58.3 40.3 22.9 41.2 49.7 - - - - - -
ResNet-50 [63] 38/44 -/260 36.3 55.3 38.6 19.3 40.0 48.8 38.0 58.6 41.4 34.4 55.1 36.7
FastViT-SA12 [14] -/- -/- - - - - - - 38.9 60.5 42.2 35.9 57.6 38.1
PVT-T [25] 23/33 221/240 36.7 56.9 38.9 22.6 38.8 50.0 36.7 59.2 39.3 35.1 56.7 37.3
EfficientFormer-L1 [43] -/32 -/196 - - - - - - 37.9 60.3 41.0 35.4 57.3 37.3
PVTv2-B0 [16] 13/24 -/- 37.2 57.2 39.5 23.1 40.4 49.7 38.2 60.5 40.7 36.2 57.8 38.6
S2AFormer-XS 13/24 166/185 37.9 58.6 40.3 22.9 41.5 49.8 38.4 60.2 41.5 35.8 57.3 38.1
SHViT-S4 [15] -/- -/- 38.8 59.8 41.1 22.0 42.4 52.7 39.0 61.2 41.9 35.9 57.9 37.9
EMO-5M [13] -/- -/- 38.9 59.8 41.0 23.8 42.2 51.7 39.3 61.7 42.4 36.4 58.4 38.7
ResNet-101 [63] 57/63 315/336 38.5 57.8 41.2 21.4 42.6 51.1 40.4 61.1 44.2 36.4 57.7 38.8
PoolFormer-S24 [30] 31/41 -/233 38.9 59.7 41.3 23.3 42.1 51.8 40.1 62.2 43.4 37.0 59.1 39.6
PoolFormer-S36 [30] 41/51 -/272 39.5 60.5 41.8 22.5 42.9 52.4 41.0 63.1 44.8 37.7 60.1 40.0
RepViT-M1.1 [31] -/- -/- - - - - - - 39.8 61.9 43.5 37.2 58.8 40.1
S2AFormer-S 17/28 178/197 40.0 60.9 42.7 24.4 43.6 52.9 41.0 62.5 45.0 37.6 59.7 40.3
PVT-S [25] 34/44 286/305 38.7 59.3 40.8 21.2 41.6 54.4 40.4 62.9 43.8 37.8 60.1 40.3
EfficientFormer-L3 [43] -/51 -/250 - - - - - - 41.4 63.9 44.7 38.1 61.0 40.4
SwiftFormer-L1 [42] -/31 -/202 - - - - - - 41.2 63.2 44.8 38.1 60.2 40.7
Swin-T [4] 38/48 248/267 41.5 62.1 44.2 25.1 44.9 55.5 42.2 64.6 46.2 39.1 61.6 42.0
EfficientFormer-L7 [43] -/101 -/378 - - - - - - 42.6 65.1 46.1 39.0 62.2 41.7
SwiftFormer-L3 [42] -/48 -/252 - - - - - - 42.7 64.4 46.7 39.1 61.7 41.8
@@
Figure 4: Qualitative results of semantic segmentation on ADE20K val dataset [79]. Upper: input validation images, Lower: results generated by S2AFormer-S-based Semantic FPN
Figure 5: Qualitative results of object detection on the COCO val2017 dataset [81]. Top: input validation images. Bottom: results generated by S2AFormer-S-based RetinaNet.
Figure 6: Qualitative results of instance segmentation on the COCO val2017 dataset [81]. Top: input validation images. Bottom: results generated by S2AFormer-S-based Mask R-CNN.
@@ IV-A Image Classification
We train the proposed S2AFormer models from scratch on ImageNet-1K [18] without using pre-trained weights or additional data. Following the strategy outlined in EdgeNeXt [44], we use a $224\times 224$ input resolution and train for 300 epochs on four NVIDIA H100 GPUs (96 GB memory). The implementation is done in PyTorch 2.5.1 with Timm 0.4.9, using AdamW [82] as the optimizer and a batch size of 1024. The initial learning rate is set to $6\times 10^{-3}$ and follows a cosine decay schedule with a 20-epoch warmup. Data augmentation includes label smoothing of 0.1, random resize cropping, horizontal flipping, RandAugment, and a multi-scale sampler. Additionally, we apply EMA [83] with a momentum factor of 0.9995.
The results on the ImageNet-1K [18] dataset, shown in Table III, highlight the significant improvements achieved by S2AFormer in image classification. Compared to state-of-the-art approaches, S2AFormer enhances classification accuracy while effectively balancing model complexity and computational efficiency.
S2AFormer-mini has 1.28M fewer parameters than SHViT-S1 [15] but outperforms it by 2.3% in accuracy. S2AFormer-T achieves 77.7% Top-1 accuracy, significantly surpassing models like SHViT-S3 [15], which has more than twice as many parameters (14.2M) and achieves only 77.4% accuracy with nearly identical MACs. Similarly, EfficientVMamba-T [70], the top-performing model in the Mamba series, falls short by 1.2% in accuracy compared to S2AFormer-T, while requiring 0.15G more MACs. S2AFormer-XS outperforms StarNet-S4 [73], which has 7.5M parameters, 1.075 GMACs, and 78.4% accuracy, while being more efficient in terms of parameters and MACs. S2AFormer-S has a slight disadvantage of 0.2% lower Top-1 accuracy compared to EdgeViT-S [12] (81.0% vs 80.8%), but it achieves this with much fewer parameters and a reduction of 0.5 GMACs, making it more lightweight. In comparison to InceptionNeXt-T [17], which has 28M parameters, 4.2 GMACs, and 82.3% accuracy, S2AFormer-M is 3.13M smaller in terms of parameter count while delivering the same Top-1 accuracy, further emphasizing its efficiency and reduced computational burden.
@@ IV-B Semantic Segmentation
We evaluate the models’ performance in semantic segmentation using the ADE20K dataset [79], a challenging benchmark for scene parsing. The dataset comprises 20,000 training images and 2,000 validation images, covering 150 detailed semantic categories. The proposed model is tested with Semantic FPN [84] as the backbone, with normalization layers frozen and pre-trained weights from ImageNet-1K [18] classification. Following standard practices [84, 85], the network is trained for 40,000 iterations with a batch size of 32, using the AdamW [82] optimizer. The initial learning rate is set to $2\times 10^{-4}$ and follows a polynomial decay schedule with a power of 0.9. During training, input images are resized and cropped to $512\times 512$. The implementation is based on the MMSegmentation [86] framework.
The comparative results shown in Table IV demonstrate that all S2AFormer variants effectively balance computational efficiency and segmentation accuracy. S2AFormer-mini, with just 6M parameters and 23 GFLOPs, achieves a competitive mIoU of 36.7%, surpassing ResNet-18 [63] and PVT-T [25] by 3.8% and 1.0%, respectively. With 7M parameters and 25 GFLOPs, S2AFormer-T improves the mIoU to 38.0%, outperforming PoolFormer-S12 [30] (16M parameters, 31 GFLOPs, 37.2% mIoU) and CASViT-XS [61] (7M parameters, 24 GFLOPs, 37.1% mIoU) by 0.8% and 0.9%, respectively, while maintaining a smaller or comparable model size. S2AFormer-XS achieves 39.2% mIoU, surpassing EfficientFormer-L1 [43] by 0.3% using fewer parameters, demonstrating its high efficiency. S2AFormer-S offers better performance than PoolFormer-S24 [30] (40.8% vs 40.3%) with fewer parameters and lower GFLOPs. Similarly, S2AFormer-M outperforms InceptionNeXt-T [17] (28M parameters, 44 GFLOPs, 43.1% mIoU) by 0.6%, while being 7.14% smaller in parameter count.
@@ IV-C Object Detection and Instance Segmentation
We assess the performance of S2AFormer on object detection and instance segmentation tasks using the COCO2017 dataset [81], which contains 118,000 training images and 5,000 validation images, along with bounding box and mask annotations across 80 categories. We assess the performance of our model with two widely used object detectors: RetinaNet [87] and Mask R-CNN [88]. To initialize the backbone, we use weights pre-trained on ImageNet. All models are trained with a batch size of 16 on four H100 GPUs, utilizing the AdamW optimizer [82] with an initial learning rate of $1\times 10^{-4}$. Following standard practices [87, 88, 89], we adopt a 1×1$\times$1 × (12 epochs) training schedule. The models are implemented using the MMDetection [89] framework.
We conducted a detailed evaluation on the COCO val2017 dataset [81], comparing backbone models based on Average Precision (AP) across various scales (small, medium, large) and tasks (detection and segmentation), along with their efficiency in terms of parameters (Para.) and FLOPs. As shown in Table LABEL:detection, S2AFormer-mini achieves an impressive 33.4% $AP$ on RetinaNet [87], surpassing models like ResNet-18 [63] and EfficientViT-M4 [11] while maintaining fewer parameters and lower GFLOPs. S2AFormer-T achieves 36.7% $AP$ for RetinaNet [87] and 37.6% $AP^{b}$ for Mask R-CNN [88], surpassing PoolFormer-S12 [30], which has 36.2% $AP$ and 37.3% in $AP^{b}$, while maintaining a reduced computational burden. Under similar parameter conditions, S2AFormer-XS (with 13/24 parameters and 166/185 GFLOPs) achieves 37.9% $AP$ for RetinaNet and 38.4% $AP^{b}$ for Mask R-CNN, outperforming PVTv2-B0 [16] (with 13/24 parameters and 37.2% $AP$ for RetinaNet and 38.2% $AP^{b}$ for Mask R-CNN) by a noticeable margin. Similarly, S2AFormer-M achieves excellent performance across benchmarks.

### v2

@@ III-C Complexity Analysis
AP AP50 AP75 APS APM APL APb $^{b}_{50} AP$$^{b}_{75}$ APm $^{m}_{50} AP$$^{m}_{75}$
MobileNetV2 [sandler2018mobilenetv2] -/- -/- 28.3 46.7 29.3 14.8 30.7 38.1 29.6 48.3 31.5 27.2 45.2 28.6
MobileNetV3 [koonce2021mobilenetv3] -/- -/- 29.9 49.3 30.8 14.9 33.3 41.1 29.2 48.6 30.3 27.1 45.5 28.2
ResNet-18 [he2016deep] 21/31 -/- 31.8 49.6 33.6 16.3 34.3 43.2 34.0 54.0 36.7 31.2 51.0 32.7
EfficientViT-M4 [liu2023efficientvit] 9/- -/- 32.7 52.2 34.1 17.6 35.3 46.0 32.8 54.4 34.5 31.0 51.2 32.2
S2AFormer-mini (Ours) 12/22 159/177 33.4 53.2 34.9 19.9 36.3 44.5 33.4 55.4 35.2 31.7 52.5 33.3
LSNet-T [wang2025lsnet] -/- -/- 34.2 54.6 35.2 17.8 37.1 48.5 35.0 57.0 37.3 32.7 53.8 34.3
PoolFormer-S12 [yu2022metaformer] 22/32 -/- 36.2 56.2 38.2 20.8 39.1 48.0 37.3 59.0 40.1 34.6 55.8 36.9
CAS-ViT-XS [zhang2024cas] 12/23 162/181 36.5 56.3 38.9 21.8 39.9 48.4 37.4 59.1 40.4 34.9 56.2 37.0
ResNet-50 [he2016deep] 38/44 -/260 36.3 55.3 38.6 19.3 40.0 48.8 38.0 58.6 41.4 34.4 55.1 36.7
FastViT-SA12 [vasu2023fastvit] -/- -/- - - - - - - 38.9 60.5 42.2 35.9 57.6 38.1
SHViT-S3 [yun2024shvit] -/- -/- 36.1 56.6 38.0 19.9 39.1 50.8 36.9 59.4 39.6 34.4 56.3 36.1
S2AFormer-T (Ours) 12/23 164/182 36.7 57.0 39.1 21.1 39.7 48.6 37.6 59.8 40.6 35.4 57.2 37.6
PVT-T [wang2021pyramid] 23/33 221/240 36.7 56.9 38.9 22.6 38.8 50.0 36.7 59.2 39.3 35.1 56.7 37.3
LSNet-S [wang2025lsnet] -/- -/- 36.7 57.2 38.6 20.0 39.7 51.8 37.4 59.9 39.8 34.8 56.8 36.6
MF-508M [chen2022mobile] 8/- 168/- 38.0 58.3 40.3 22.9 41.2 49.7 - - - - - -
EfficientFormer-L1 [li2022efficientformer] -/32 -/196 - - - - - - 37.9 60.3 41.0 35.4 57.3 37.3
PVTv2-B0 [wang2021pvtv2] 13/24 -/- 37.2 57.2 39.5 23.1 40.4 49.7 38.2 60.5 40.7 36.2 57.8 38.6
S2AFormer-XS (Ours) 13/24 166/185 37.9 58.6 40.3 22.9 41.5 49.8 38.4 60.2 41.5 35.8 57.3 38.1
SHViT-S4 [yun2024shvit] -/- -/- 38.8 59.8 41.1 22.0 42.4 52.7 39.0 61.2 41.9 35.9 57.9 37.9
EMO-5M [zhang2023rethinking] -/- -/- 38.9 59.8 41.0 23.8 42.2 51.7 39.3 61.7 42.4 36.4 58.4 38.7
ResNet-101 [he2016deep] 57/63 315/336 38.5 57.8 41.2 21.4 42.6 51.1 40.4 61.1 44.2 36.4 57.7 38.8
EfficientViM-M4 [lee2025efficientvim] -/- -/- 38.8 59.6 41.1 22.1 42.4 52.8 39.3 60.2 42.5 35.8 57.1 37.4
PoolFormer-S24 [yu2022metaformer] 31/41 -/233 38.9 59.7 41.3 23.3 42.1 51.8 40.1 62.2 43.4 37.0 59.1 39.6
PoolFormer-S36 [yu2022metaformer] 41/51 -/272 39.5 60.5 41.8 22.5 42.9 52.4 41.0 63.1 44.8 37.7 60.1 40.0
RepViT-M1.1 [wang2024repvit] -/- -/- - - - - - - 39.8 61.9 43.5 37.2 58.8 40.1
S2AFormer-S (Ours) 17/28 178/197 40.0 60.9 42.7 24.4 43.6 52.9 41.0 62.5 45.0 37.6 59.7 40.3
PVT-S [wang2021pyramid] 34/44 286/305 38.7 59.3 40.8 21.2 41.6 54.4 40.4 62.9 43.8 37.8 60.1 40.3
EfficientFormer-L3 [li2022efficientformer] -/51 -/250 - - - - - - 41.4 63.9 44.7 38.1 61.0 40.4
SwiftFormer-L1 [shaker2023swiftformer] -/31 -/202 - - - - - - 41.2 63.2 44.8 38.1 60.2 40.7
Swin-T [liu2021swin] 38/48 248/267 41.5 62.1 44.2 25.1 44.9 55.5 42.2 64.6 46.2 39.1 61.6 42.0
EfficientFormer-L7 [li2022efficientformer] -/101 -/378 - - - - - - 42.6 65.1 46.1 39.0 62.2 41.7
SwiftFormer-L3 [shaker2023swiftformer] -/48 -/252 - - - - - - 42.7 64.4 46.7 39.1 61.7 41.8
@@ IV-A Image Classification
We train the proposed S2AFormer models from scratch on ImageNet-1K [deng2009imagenet] without using pre-trained weights or additional data. Following the strategy outlined in EdgeNeXt [maaz2022edgenext], we use a $224\times 224$ input resolution and train for 300 epochs on four NVIDIA H100 GPUs (96 GB memory). The implementation is done in PyTorch 2.5.1 with Timm 0.4.9, using AdamW [loshchilov2017decoupled] as the optimizer and a batch size of 1024. The initial learning rate is set to $6\times 10^{-3}$ and follows a cosine decay schedule with a 20-epoch warmup. Data augmentation includes label smoothing of 0.1, random resize cropping, horizontal flipping, RandAugment, and a multi-scale sampler. Additionally, we apply EMA [polyak1992acceleration] with a momentum factor of 0.9995.
The results on the ImageNet-1K [deng2009imagenet] dataset, shown in Table IVb, highlight the significant improvements achieved by S2AFormer in image classification. Compared to state-of-the-art approaches, S2AFormer enhances classification accuracy while effectively balancing model complexity and computational efficiency. S2AFormer-mini has 1.28M fewer parameters than that of SHViT-S1 [yun2024shvit] but outperforms it by 2.3% in accuracy. S2AFormer-T achieves 77.7% Top-1 accuracy, significantly surpassing models like SHViT-S3 [yun2024shvit], which has more than twice as many parameters (14.2M) and achieves only 77.4% accuracy with nearly identical MACs. Similarly, EfficientVMamba-T [pei2024efficientvmamba], the top-performing model in the Mamba series, falls short by 1.2% in accuracy compared to S2AFormer-T, while requiring 0.15G more MACs. S2AFormer-XS outperforms StarNet-S4 [ma2024rewrite], which has 7.5M parameters, 1.075 GMACs, and 78.4% accuracy, while being more efficient in terms of parameters and MACs. S2AFormer-S has a slight disadvantage of 0.2% lower Top-1 accuracy compared to EdgeViT-S [chen2022edgevit] (81.0% vs 80.8%), but it achieves this with much fewer parameters and a reduction of 0.5 GMACs, making it more lightweight. In comparison to InceptionNeXt-T [yu2024inceptionnext], which has 28M parameters, 4.2 GMACs, and 82.3% accuracy, S2AFormer-M is 3.13M smaller in terms of parameter count while delivering the same Top-1 accuracy, further emphasizing its efficiency and reduced computational burden.
@@ IV-B Semantic Segmentation
We evaluate the models’ performance in semantic segmentation using the ADE20K dataset [zhou2017scene], a challenging benchmark for scene parsing. The dataset comprises 20,000 training images and 2,000 validation images, covering 150 detailed semantic categories. The proposed model is tested with Semantic FPN [kirillov2019panoptic] as the backbone, with normalization layers frozen and pre-trained weights from ImageNet-1K [deng2009imagenet] classification. Following standard practices [kirillov2019panoptic, chen2017deeplab], the network is trained for 40,000 iterations with a batch size of 32, using the AdamW [loshchilov2017decoupled] optimizer. The initial learning rate is set to $2\times 10^{-4}$ and follows a polynomial decay schedule with a power of 0.9. During training, input images are resized and cropped to $512\times 512$. The implementation is based on the MMSegmentation [mmseg2020] framework.
The comparative results shown in Table V demonstrate that all S2AFormer variants effectively balance computational efficiency and segmentation accuracy. S2AFormer-mini, with just 6M parameters and 23 GFLOPs, achieves a competitive mIoU of 36.7%, surpassing ResNet-18 [he2016deep] and PVT-T [wang2021pyramid] by 3.8% and 1.0%, respectively. With 7M parameters and 25 GFLOPs, S2AFormer-T improves the mIoU to 38.0%, outperforming PoolFormer-S12 [yu2022metaformer] (16M parameters, 31 GFLOPs, 37.2% mIoU) and CASViT-XS [zhang2024cas] (7M parameters, 24 GFLOPs, 37.1% mIoU) by 0.8% and 0.9%, respectively, while maintaining a smaller or comparable model size. S2AFormer-XS achieves 39.2% mIoU, surpassing EfficientFormer-L1 [li2022efficientformer] by 0.3% using fewer parameters, demonstrating its high efficiency. S2AFormer-S offers better performance than PoolFormer-S24 [yu2022metaformer] (40.8% vs 40.3%) with fewer parameters and lower GFLOPs. Similarly, S2AFormer-M outperforms InceptionNeXt-T [yu2024inceptionnext] (28M parameters, 44 GFLOPs, 43.1% mIoU) by 0.6%, while being 7.14% smaller in parameter count.
@@ IV-C Object Detection and Instance Segmentation
We assess the performance of S2AFormer on object detection and instance segmentation tasks using the COCO2017 dataset [lin2014microsoft], which contains 118,000 training images and 5,000 validation images, along with bounding box and mask annotations across 80 categories. We assess the performance of our model with two widely used object detectors: RetinaNet [ross2017focal] and Mask R-CNN [he2017mask]. To initialize the backbone, we use weights pre-trained on ImageNet. All models are trained with a batch size of 16 on four H100 GPUs, utilizing the AdamW optimizer [loshchilov2017decoupled] with an initial learning rate of $1\times 10^{-4}$. Following standard practices [ross2017focal, he2017mask, chen2019mmdetection], we adopt a $1\times$ (12 epochs) training schedule. The models are implemented using the MMDetection [chen2019mmdetection] framework.
We conducted a detailed evaluation on the COCO val2017 dataset [lin2014microsoft], comparing backbone models based on Average Precision (AP) across various scales (small, medium, large) and tasks (detection and segmentation), along with their efficiency in terms of parameters (Para.) and FLOPs. As shown in Table LABEL:detection, S2AFormer-mini achieves an impressive 33.4% $AP$ on RetinaNet [ross2017focal], surpassing models like ResNet-18 [he2016deep] and EfficientViT-M4 [liu2023efficientvit] while maintaining fewer parameters and lower GFLOPs. S2AFormer-T achieves 36.7% $AP$ for RetinaNet [ross2017focal] and 37.6% $AP^{b}$ for Mask R-CNN [he2017mask], surpassing PoolFormer-S12 [yu2022metaformer], which has 36.2% $AP$ and 37.3% in $AP^{b}$, while maintaining a reduced computational burden. Under similar parameter conditions, S2AFormer-XS (with 13/24 parameters and 166/185 GFLOPs) achieves 37.9% $AP$ for RetinaNet and 38.4% $AP^{b}$ for Mask R-CNN, outperforming PVTv2-B0 [wang2021pvtv2] (with 13/24 parameters and 37.2% $AP$ for RetinaNet and 38.2% $AP^{b}$ for Mask R-CNN) by a noticeable margin. Similarly, S2AFormer-M achieves excellent performance across benchmarks.
Figure 6: Qualitative results of semantic segmentation on ADE20K val dataset [zhou2017scene]. Upper: input validation images, Lower: results generated by S2AFormer-S-based Semantic FPN [kirillov2019panoptic].
Figure 7: Qualitative results of object detection on the COCO val2017 dataset [lin2014microsoft]. Top: input validation images. Bottom: results generated by S2AFormer-S-based RetinaNet [ross2017focal].
Figure 8: Qualitative results of instance segmentation on the COCO val2017 dataset [lin2014microsoft]. Top: input validation images. Bottom: results generated by S2AFormer-S-based Mask R-CNN [he2017mask].